Vic has 14 coins in nickels and quarters. The value of his coins is $1.70. How many quarters does he have?
Answer by mananth(16946)

let quarters be x
nickels will be 14-x
...
25x+5(14-x)=170
25x+70-5x=170
-70 -70
20x=100
/20
x= 5 quarters
...
m.ananth@hotmail.ca